# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ Abstract.
We develop a unified relative entropy framework for macroscopic limits of kinetic equations with Riesz-type interactions and Fokker–Planck relaxation. The method combines entropy dissipation, Fisher-information control, and modulated interaction energies into a robust stability theory that yields both strong and weak convergence results. For the strong convergence, we establish quantitative relative entropy estimates toward macroscopic limits under well-prepared data, extending the scope of the method to settings where nonlocal forces and singular scalings play a decisive role. For the weak convergence, we prove that quantitative convergence propagates in bounded Lipschitz topologies, even when the initial relative entropy diverges with respect to the singular scaling parameter. This dual perspective shows that relative entropy provides not only a tool for strong convergence, but also a new mechanism to handle mildly prepared initial states. We establish quantitative convergence toward three prototypical limits: the diffusive limit leading to a drift-diffusion equation, the high-field limit yielding the aggregation equation in the repulsive regime, and the strong magnetic field limit producing a generalized surface quasi-geostrophic equation. The analysis highlights the unifying role of relative entropy in connecting microscopic dissipation with both strong and weak macroscopic convergence.
@@ 1. Introduction
A key challenge in this program lies in handling singular scalings, nonlocal interactions, or strong external fields. In these regimes, the usual compactness or moment methods do not provide sufficient control, and more refined tools such as entropy methods, Wasserstein stability estimates, or modulated energies have been developed, e.g. [5, 6, 14, 15, 16, 18, 21, 24, 27, 28, 37, 39, 43, 56, 57, 59]. Establishing quantitative convergence results in such singular limits is particularly important: they not only provide error bounds for the approximation of kinetic dynamics by macroscopic models but also clarify the regimes of validity and the robustness of effective descriptions under weakly prepared data.
@@
  • gSQG limit. In the strong magnetic field regime ${\rm A}=\varepsilon$, ${\rm B}=\tau=1$, the additional fast rotational term constrains the dynamics to two dimensions, leading to a generalized surface quasi-geostrophic (gSQG) type equation.
Beyond these specific regimes, our approach relies on a unified relative entropy framework that applies to all three limits. This method provides not only qualitative convergence but also quantitative bounds. In particular, we establish stability estimates in strong topologies and extend the analysis to weak metrics by employing the bounded Lipschitz distance, which yields additional quantitative convergence results.
@@ 1.1. Functional frameworks
  • Distance functionals, such as the bounded Lipschitz (BL) distance, which allow us to measure weak convergence of distributions and to formulate stability estimates under less restrictive assumptions on the initial data.
@@ 1.1.1. Relative entropy
It quantifies the deviation of $f$ from the local equilibrium $M_{\rho,u}$. The functional (1.2) is nonnegative, convex in $f$, and vanishes if and only if $f=M_{\rho,u}$. Equivalently, formula (1.2) reveals that for prescribed $(\rho,u)$, the local Maxwellian $M_{\rho,u}$ is the unique minimizer of the variational problem:
This characterization can be justified in several ways:
gives the identity
$\log g(\xi)+1+\frac{1}{2}|\xi|^{2}+\lambda_{0}+\lambda\cdot\xi=0,$
which implies
Completing the square in the exponent,
so that
$g(\xi)=C_{0}e^{\frac{1}{2}|\lambda|^{2}}e^{-\frac{1}{2}|\xi+\lambda|^{2}}.$
$C_{0}e^{\frac{1}{2}|\lambda|^{2}}=\frac{\rho}{(2\pi)^{\frac{d}{2}}},$
$g(\xi)=M_{\rho,u}(\xi),$
which shows that the Maxwellian is the unique minimizer of (1.3).
@@ 1.1.2. Energy and modulated functionals
which accounts for the nonlocal forces induced by the Riesz kernel. The operator $(-\Delta)^{-\alpha}$ can be expressed in terms of convolution with the Riesz potential
@@
For $\alpha=1$, we instead use the classical Coulomb potential representation, equivalently
$\mathscr{P}[\rho]=\int_{\mathbb{R}^{d}}|\nabla_{x}\phi|^{2}\,dx\quad\mbox{with }-\Delta_{x}\phi=\rho.$
@@
$\mathscr{P}[\rho_{f}|\rho]=\frac{1}{2}\int_{\mathbb{T}^{d}}(\rho_{f}-\rho)(-\Delta)^{-\alpha}(\rho_{f}-\rho)\,dx=\frac{1}{2}\|\rho_{f}-\rho\|_{\dot{H}^{-\alpha}(\mathbb{T}^{d})}^{2}$
@@ 1.1.3. Bounded Lipschitz distance
Besides entropy-based functionals, we also employ weak metrics to quantify convergence of distributions. For a metric space $(X,{\rm d}_{X})$ and finite Borel measures $\mu,\nu$ on $X$, the BL distance is defined by
@@
Beyond these structural advantages, the BL distance plays a crucial role in our quantitative estimates. Through Lemma 3.2 below, the dissipation terms appearing in the relative entropy inequality can be converted into control of time-integrated BL distances between $f^{\varepsilon}$ and the corresponding local Maxwellians. This mechanism allows us to complement the relative entropy framework with quantitative weak stability bounds. In particular, it enables us to treat classes of initial data that are not strongly aligned with the limiting profile. The precise notions of “well-prepared” and “mildly well-prepared” data, and the corresponding stability results, will be introduced in the next subsection on main results.
Before closing this part, we introduce several simplified notations that will be used throughout this paper. For functions $f(x,v)$ and $g(x)$, when there is no risk of confusion, we will use $\left\|f\right\|_{L^{p}}$ and $\left\|g\right\|_{L^{p}}$ to denote the usual $L^{p}(\mathbb{R}^{d}\times\mathbb{R}^{d})$-norm and $L^{p}(\mathbb{R}^{d})$-norm, respectively. We denote by $\mathcal{M}(\mathbb{R}^{n})$ the space of finite signed Radon measures on $\mathbb{R}^{n}$, and by $\mathcal{P}_{p}(\mathbb{R}^{n})$, $p\in[1,\infty]$, the space of probability measures on $\mathbb{R}^{n}$ with finite $p$-th moment. We also use a velocity-weighted $L^{1}$-norm defined by
@@ 1.2.1. Diffusive limit
This formal derivation indicates that, in the limit $\varepsilon\to 0$, the macroscopic density $\rho$ evolves according to the aggregation–diffusion equation (1.7). We now make this rigorous: the next result shows that weak entropy solutions of the kinetic equation (1.5) converge, under suitable assumptions on the initial data, to the regular solution $\rho$ of (1.7), and that quantitative stability estimates can be obtained in this setting.
@@
If the initial data is mildly well-prepared in the sense that
${\rm d}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})\to 0\quad\mbox{and}\quad\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\right)\to 0\quad\mbox{as }\varepsilon\to 0,$
then we have
$f^{\varepsilon}$ $\overset{*}{\rightharpoonup}M_{\rho,0}\quad\mbox{in }L^{2}(0,T;\mathcal{M}(\mathbb{R}^{d}\times\mathbb{R}^{d})),$
$\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\xi f^{\varepsilon}\,d\xi$ $\overset{*}{\rightharpoonup}-\rho\left(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho\right)\quad\mbox{in }\mathcal{M}((0,T)\times\mathbb{R}^{d}).$
Moreover, we obtain the quantitative bounds
$\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\,dt$ $\leq C{\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})+C\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\right)+C\varepsilon^{2},$
$\sup_{0\leq t\leq T}{\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}}(t),\rho(t))$ $\leq C{\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})+C\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\right)+C\varepsilon^{4}.$
@@
The notion of mildly well-prepared initial data can be interpreted as an intermediate regime between the classical well-prepared and the fully ill-prepared settings. In contrast to the standard well-prepared case, we do not require the modulated energy $\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)$ to vanish as $\varepsilon\to 0$. Our framework permits substantially larger deviations of the kinetic initial state from equilibrium, and in fact accommodates situations where this quantity diverges, with growth up to order
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\sim\frac{1}{\varepsilon^{2-}}\quad\mbox{as }\varepsilon\to 0.$
To illustrate this, consider the initial data
$f^{\varepsilon}_{0}(x,\xi)=\frac{\rho_{0}(x)}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{|u_{\varepsilon}(x)-\xi|^{2}}{2}\right),\quad u_{\varepsilon}(x)=\frac{v(x)}{\varepsilon^{1-\frac{\delta}{2}}},$
where $v\in L^{2}(\rho_{0}\,dx)$ and $\delta\in(0,2)$, and assume
$\rho_{0}\in\mathcal{P}\cap\dot{H}^{-\alpha}(\mathbb{R}^{d}).$
Then, we obtain
$\rho_{f^{\varepsilon}_{0}}=\int_{\mathbb{R}^{d}}f^{\varepsilon}_{0}\,d\xi=\rho_{0}\quad\mbox{and}\quad\mathscr{K}[f^{\varepsilon}_{0}]+\mathscr{P}[\rho_{f^{\varepsilon}_{0}}]=\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{0}|u_{\varepsilon}|^{2}\,dx+\mathscr{P}[\rho_{0}]+\frac{d}{2}.$
In particular,
$\varepsilon^{2}\left(\mathscr{K}[f^{\varepsilon}_{0}]+\mathscr{P}[\rho_{f^{\varepsilon}_{0}}]\right)=\varepsilon^{2}\left(\mathscr{P}[\rho_{0}]+\frac{d}{2}\right)+\frac{\varepsilon^{\delta}}{2}\int_{\mathbb{R}^{d}}\rho_{0}|v|^{2}\,dx<\infty$
uniformly in $\varepsilon\in(0,1)$. Since $\rho_{f^{\varepsilon}_{0}}=\rho_{0}$, we further have
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)=\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)=\mathscr{F}[f^{\varepsilon}_{0}]-\int_{\mathbb{R}^{d}}\rho_{0}\log\rho_{0}\,dx=\frac{1}{2}\int_{\mathbb{R}^{d}}\rho_{0}|u_{\varepsilon}|^{2}\,dx,$
and therefore
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)=\frac{1}{2\varepsilon^{2-\delta}}\int_{\mathbb{R}^{d}}\rho_{0}|v|^{2}\,dx\sim\frac{1}{\varepsilon^{2-\delta}}$
for any $\delta\in(0,2)$.
This shows that our convergence result remains valid well beyond the classical well-prepared regime, and encompasses a range of ill-prepared configurations with large initial modulated energies.
${\rm d}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})\to 0,$
which enforces that the macroscopic density is correctly aligned with the limiting initial data. This condition is not specific to our theorem: it is in fact essential in any framework aiming at quantitative convergence estimates, since without such consistency at the level of densities, one cannot close stability inequalities. If this condition is removed, only qualitative convergence can be expected, typically via compactness arguments (see [35] for example). Thus, the mildly well-prepared regime should be seen as allowing highly unbalanced kinetic initial states in terms of energy, while still requiring the local density to be prepared, in order to retain access to quantitative stability bounds.
Remark 1.3.
If we impose a stronger well-prepared condition on the initial data, namely
${\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)=O(\varepsilon^{2}),$
then the convergence rate improves to
$\sup_{0\leq t\leq T}{\rm d}_{\rm BL}(\rho_{f^{\varepsilon}}(t),\rho(t))\leq C\varepsilon^{2}.$
This indicates that in the regime of genuinely well-prepared initial data, convergence is not only guaranteed but also improves beyond the previously known $O(\varepsilon)$ rate. In fact, our result provides an $O(\varepsilon^{2})$ bound, which corresponds to the natural diffusive scaling: under the kinetic rescaling leading to the diffusion equation, the characteristic error terms are of order $\varepsilon^{2}$, and our estimate matches this scaling. To the best of our knowledge, such a rate has not been established in earlier works.
We stress that this result complements the previous remark: while the mildly well-prepared regime allows for large initial deviations in energy (and thus only ensures weaker convergence), the stronger alignment of the initial modulated energies considered here produces a significantly improved quantitative bound. This highlights how the degree of well-preparedness of the initial data directly determines the attainable convergence rate within the relative entropy method.
Remark 1.4.
In the case of mildly well-prepared initial data, if the initial total energy $\mathscr{F}[f^{\varepsilon}_{0}]+\mathscr{P}[\rho_{f^{\varepsilon}_{0}}]$ is uniformly bounded in $\varepsilon$, we can also establish the convergence estimates
$\int_{\mathbb{R}^{d}}f^{\varepsilon}\,d\xi$ $\rightharpoonup\rho\quad\mbox{in }L^{1}((0,T)\times\mathbb{R}^{d}),$
$\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\xi f^{\varepsilon}\,d\xi$ $\rightharpoonup-\rho\left(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho\right)\quad\mbox{in }L^{1}((0,T)\times\mathbb{R}^{d})$
under the additional assumption of the uniform spatial moment bound
$\sup_{0<\varepsilon<1}\||x|f^{\varepsilon}_{0}\|_{L^{1}}<\infty.$
Indeed, the uniform bound on the total energy implies that the entropy term $\rho_{f^{\varepsilon}}\log\rho_{f^{\varepsilon}}$ is uniformly bounded in $L^{\infty}(0,T;L^{1}(\mathbb{R}^{d}))$, which gives the equiintegrability of $\rho_{f^{\varepsilon}}$ in $L^{\infty}(0,T;L^{1}(\mathbb{R}^{d}))$. Then, by the Dunford–Pettis theorem, combined with tightness provided by the moment and kinetic energy bounds, we deduce that the density and the flux are relatively compact in $L^{1}$.
In the Coulombic potential case, one can further obtain the strong convergences
$f^{\varepsilon}$ $\to M_{\rho,0}\quad\mbox{in }L^{1}(0,T;L^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d})),$
$\int_{\mathbb{R}^{d}}f^{\varepsilon}\,d\xi$ $\to\rho\quad\mbox{in }L^{2}((0,T)\times\mathbb{R}^{d}),$
by employing additional compactness arguments such as the Kolmogorov–Riesz–Fréchet theorem; see [28, 35].
@@
The proof strategy used for the Manev potential in [25] can in fact be extended to cover the whole range $\alpha\in[\tfrac{1}{2},1)$, although rigorous results in this direction are not yet available in the literature. By contrast, for $\alpha\in(0,\tfrac{1}{2})$, the global-in-time existence of renormalized or weak solutions, to the best of our knowledge, remains open and appears to be highly challenging. However, for small and smooth initial data, global solutions can be constructed; see, for example, [26]. For the existence of solutions to the aggregation–diffusion equation (1.7) with the regularity required in Theorem 1.1, we provide a detailed discussion in Appendix B.

### v2

@@ Abstract.
We develop a unified relative entropy framework for macroscopic limits of kinetic equations with Riesz-type interactions and Fokker–Planck relaxation. Our analysis covers three prototypical singular regimes: the diffusive limit leading to a drift-diffusion equation, the high-field limit yielding the aggregation equation in the repulsive regime, and the strong magnetic field limit producing a generalized surface quasi-geostrophic equation. The method combines entropy dissipation, Fisher-information control, and modulated interaction energies into a robust stability theory yielding both strong and weak convergence results. For the strong convergence, we establish quantitative relative entropy estimates toward macroscopic limits under well-prepared initial data, extending the scope of the method to settings where nonlocal forces and singular scalings play a decisive role. For the weak convergence, our approach captures three complementary phenomena: in the diffusive regime, it yields sharper quantitative estimates in weak topologies consistent with the formally optimal scaling; in the high-field regime, it propagates bounded Lipschitz stability for a class of mildly prepared initial data, even when the relative entropy diverges with respect to the singular scaling parameter; and in the strong magnetic field regime, it provides quantitative weak estimates, including bounded Lipschitz control of the rescaled momentum and negative Sobolev control of the density. This broader perspective shows that relative entropy provides not only a tool for strong convergence, but also a mechanism for treating low-regularity and mildly prepared regimes. The analysis highlights the unifying role of relative entropy in connecting microscopic dissipation with both strong and weak macroscopic convergence.
@@ 1. Introduction
A key challenge in this program lies in handling singular scalings, nonlocal interactions, or strong external fields. In these regimes, the usual compactness or moment methods do not provide sufficient control, and more refined tools such as entropy methods, Wasserstein stability estimates, or modulated energies have been developed, e.g. [5, 6, 17, 18, 14, 16, 30, 25, 28, 29, 39, 41, 45, 59, 60, 62]. Establishing quantitative convergence results in such singular limits is particularly important: they not only provide error bounds for the approximation of kinetic dynamics by macroscopic models but also clarify the regimes of validity and the robustness of effective descriptions in low-regularity or partially prepared settings.
@@
  • gSQG limit. In the strong magnetic field regime ${\rm A}=\varepsilon$, ${\rm B}=\tau=1$, the kinetic equation contains an additional fast rotational transport term of the form
    $\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f,$
    which represents the cyclotronic motion generated by a constant external magnetic field. This gyroscopic force induces rapid angular averaging in velocity space and effectively constrains the dynamics to two dimensions. In combination with the Riesz interaction, the resulting macroscopic evolution is a generalized surface quasi-geostrophic (gSQG) equation with active scalar velocity $\nabla_{x}^{\perp}(-\Delta_{x})^{-\alpha}\rho$.
Beyond these specific regimes, our approach relies on a unified relative entropy framework that applies to all three limits. This method provides not only qualitative convergence but also quantitative bounds. In particular, we establish stability estimates in strong topologies and, in appropriate regimes, obtain additional quantitative convergence results in weak topologies, including bounded Lipschitz distances and negative Sobolev estimates.
@@ 1.1. Functional frameworks
  • Distance functionals and weak topologies, including the bounded Lipschitz (BL) distance and negative Sobolev norms, which allow us to quantify convergence in weaker topologies and to obtain stability estimates adapted to different singular regimes.
@@ 1.1.1. Relative entropy
It quantifies the deviation of $f$ from the local equilibrium $M_{\rho,u}$. The functional (1.2) is nonnegative, convex in $f$, and vanishes if and only if $f=M_{\rho,u}$. Moreover, it is closely related to the variational characterization of Maxwellians under prescribed mass and momentum constraints; see Section 2 below.
In the present work, relative entropy serves as the basic microscopic stability functional. Combined with the interaction part of the modulated energy, it allows us to measure simultaneously the deviation from local Maxwellians and the mismatch of the associated density variable. This two-level structure is one of the key ingredients of our unified approach to the three singular limits considered in this paper. More specifically, relative entropy admits a natural decomposition into a microscopic part, measuring the deviation from the Maxwellian with the same local mass and momentum as $f$, and a macroscopic part, measuring the mismatch between these moments and the target state $(\rho,u)$. This structure will be made precise in Section 2; it is one of the key mechanisms that allows the relative entropy method to connect microscopic dissipation with macroscopic stability.
@@ 1.1.2. Energy and modulated functionals
which accounts for the nonlocal forces induced by the Riesz kernel. The operator $(-\Delta_{x})^{-\alpha}$ can be expressed in terms of convolution with the Riesz potential
@@
$\mathscr{P}[\rho_{f}|\rho]=\frac{1}{2}\int_{\mathbb{T}^{d}}(\rho_{f}-\rho)(-\Delta_{x})^{-\alpha}(\rho_{f}-\rho)\,dx=\frac{1}{2}\|\rho_{f}-\rho\|_{\dot{H}^{-\alpha}(\mathbb{T}^{d})}^{2}$
@@ 1.1.3. Bounded Lipschitz distance
Besides entropy-based functionals, we also employ weaker topologies to quantify convergence of distributions. Depending on the scaling regime, these are measured either by bounded Lipschitz distances or by negative Sobolev norms. In particular, while negative Sobolev spaces are useful for capturing quantitative convergence in the diffusive regime, the bounded Lipschitz distance is especially well-suited for weak stability estimates at the level of measures.
For a metric space $(X,{\rm d}_{X})$ and finite Borel measures $\mu,\nu$ on $X$, the BL distance is defined by
@@
Beyond these structural advantages, the BL distance plays a crucial role in our quantitative estimates. Through Lemma 3.2 below, the dissipation terms appearing in the relative entropy inequality can be converted into control of time-integrated BL distances between $f^{\varepsilon}$ and the corresponding local Maxwellians. This mechanism allows us to complement the relative entropy framework with quantitative weak stability bounds. It is particularly useful for the high-field limit, where weak convergence estimates are established under a mildly well-prepared assumption on the initial data.
Before closing this part, we introduce several simplified notations that will be used throughout this paper. For functions $f(x,v)$ and $g(x)$, when there is no risk of confusion, we will use $\|f\|_{L^{p}}$ and $\|g\|_{L^{p}}$ to denote the usual $L^{p}(\mathbb{R}^{d}\times\mathbb{R}^{d})$-norm and $L^{p}(\mathbb{R}^{d})$-norm, respectively. We denote by $\mathcal{M}(\mathbb{R}^{n})$ the space of finite signed Radon measures on $\mathbb{R}^{n}$, and by $\mathcal{P}_{p}(\mathbb{R}^{n})$, $p\in[1,\infty]$, the space of probability measures on $\mathbb{R}^{n}$ with finite $p$-th moment. We also use a velocity-weighted $L^{1}$-norm defined by
@@ 1.2.1. Diffusive limit
This formal derivation indicates that, in the limit $\varepsilon\to 0$, the macroscopic density $\rho$ evolves according to the aggregation–diffusion equation (1.6). We now make this rigorous. More precisely, under suitable assumptions on the initial data, weak entropy solutions of the kinetic equation (1.4) converge to the regular solution $\rho$ of (1.6).
Our first result provides quantitative convergence estimates in the natural modulated energy topology, yielding strong control of both the kinetic density and the momentum variable. In addition, by exploiting the parabolic structure of the limiting equation, we also derive a sharper estimate for the density in a weaker topology of negative Sobolev type. This low-regularity estimate is particularly useful in the diffusive regime: although the topology is weaker, it is well adapted to the kinetic-to-macroscopic passage and captures the optimal-order convergence rate predicted by the formal expansion.
@@
If we further assume that $\rho\in L^{\infty}(0,T;H^{\beta+1}(\mathbb{R}^{d}))$ for some $\beta>\frac{d}{2}+1$, then the density error also satisfies the following quantitative estimate in the negative Sobolev space $H^{-\beta}(\mathbb{R}^{d})$:
$\|\rho_{f^{\varepsilon}}-\rho\|_{L^{2}(0,T;H^{-\beta})}\leq C\left(\|\rho_{f_{0}^{\varepsilon}}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)+\varepsilon^{2}\right).$
@@
Formal computations in Appendix B.1 show that
$f^{\varepsilon}=M_{\rho,0}+\varepsilon f_{1}+O(\varepsilon^{2})\quad\text{with }\int_{\mathbb{R}^{d}}f_{1}\,d\xi=0.$
Accordingly, one expects the convergence rates
$f^{\varepsilon}-M_{\rho,0}=O(\varepsilon),\quad\rho_{f^{\varepsilon}}-\rho=O(\varepsilon^{2}),\quad\text{and}\quad\rho_{f^{\varepsilon}}\frac{u_{f^{\varepsilon}}}{\varepsilon}+\rho(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho)=O(\varepsilon)$
to be essentially optimal.
Moreover, introducing the effective velocity
${\rm u}_{\varepsilon}:=-\varepsilon(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho),$
the first-order corrector is exactly captured by the linearization of the local Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ around $M_{\rho,0}$. Indeed, one formally obtains the improved approximation
$f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}=O(\varepsilon^{2}).$
Thus, while $M_{\rho,0}$ approximates the microscopic distribution only at order $O(\varepsilon)$, the shifted Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ incorporates the leading drift correction and provides a second-order approximation. This refined convergence can also be justified rigorously in a weaker negative Sobolev topology, see Remark 4.1 for more details below.
In particular, the estimate in the weaker topology of negative Sobolev spaces is consistent with the diffusive scaling and provides a natural framework in which this optimal-order behavior can be quantified.
@@
The proof strategy used for the Manev potential in [26] can in fact be extended to cover the whole range $\alpha\in[\tfrac{1}{2},1)$, although rigorous results in this direction are not yet available in the literature. By contrast, for $\alpha\in(0,\tfrac{1}{2})$, the global-in-time existence of renormalized or weak solutions, to the best of our knowledge, remains open and appears to be highly challenging. However, for small and smooth initial data, global solutions can be constructed; see, for example, [21, 27, 23]. For the existence of solutions to the aggregation–diffusion equation (1.6) with the regularity required in Theorem 1.1, we provide a detailed discussion in Appendix C.